OpenStudy (anonymous):

The point-slope form of the equation of a line that passes through points (8, 4) and (0, 2) is y – 4 = (x – 8). What is the slope-intercept form of the equation for this line? y = x – 12 y = x – 4 y = x + 2 y = x + 6

OpenStudy (teddyiswatshecallsme):

You start out with y - 4 = (x - 8) keep in mind eight is a negative eight. so you add 4 to both sides to isolate y and then that gives you y = x - 4